We have 4 positive numbers whose paired products are 2, 3, 4, 5, 6 and x. Find the 4 numbers and x.
1
Expert's answer
2020-04-08T10:32:04-0400

Let's denote the four numbers we are looking for by a, b, c and d respectively.

We do not know which of the pairs of numbers give the corresponding products. Therefore only the general formulas are valid, in which we denote the products by unknown p1, p2, p3, p4, p5 and X (the equations are enumerated for convenience)

ab=p1 (1),

ac=p2 (2),

ad=p3 (3),

bc=p4 (4),

bd=p5 (5),

cd=X (6),

where we can assign the expression for the product X arbitrarily.

From (1) we get b=p1/a (7),

from (2) we get c=p2/a (8),

from (13) we get d=p3/a (9).

From (4), (7) and (8) we get a2=p1p2/p4 (10),

from (5), (7) and (9) we get a2=p1p3/p5 (11),

from (6), (8) and (9) we get X=p2p3/a2 (12).


From (10) and (11) we can see that p2/p4=p3/p5 (13).

If we take into account the numerical values of the products (2, 3, 4, 5, 6), then the equality (13) is possible only in two ways:

Case 1: p2=2, p4=3, p3=4 and p5=6

Case 2: p2=2, p4=4, p3=3 and p5=6

Other permutations are equivalent.


Consider case 1: p2=2, p4=3, p3=4 and p5=6. In this case p1=5.

From (10) we get a2=p1p2/p4,

a2=5*2/3.

a="\sqrt{\frac{10}{3}}" because according to the condition of the problem the numbers are positive.

From (7) we get b=p1/a=5/"\sqrt{\frac{10}{3}}" =5"\sqrt{\frac{3}{10}}" ="\sqrt{\frac{15}{2}}" ,

from (8) we get c=p2/a=2/"\sqrt{\frac{10}{3}}" =2"\sqrt{\frac{3}{10}}" ="\sqrt{\frac{6}{5}}" ,

from (9) we get d=p3/a=4/"\sqrt{\frac{10}{3}}" =4"\sqrt{\frac{3}{10}}" ="\sqrt{\frac{24}{5}}" .

And from (12) we get X=p2p3/a2=2*4/(10/3)=12/5.


Consider case 2: p2=2, p4=4, p3=3 and p5=6.

In this case p1=5 again.

From (10) we get a2=p1p2/p4,

a2=5*2/4,

a="\sqrt{\frac{5}{2}}" because according to the condition of the task the numbers are positive

From (7) we get b=p1/a=5/"\sqrt{\frac{5}{2}}" =5"\sqrt{\frac{2}{5}}" ="\sqrt{10}" ,

from (8) we get c=p2/a=2/"\sqrt{\frac{5}{2}}" =2"\sqrt{\frac{2}{5}}" ="\sqrt{\frac{8}{5}}" ,

from (9) we get d=p3/a=3/"\sqrt{\frac{5}{2}}" =3"\sqrt{\frac{2}{5}}" ="\sqrt{\frac{18}{5}}" .

And from (12) we get X=p2p3/a2=3*6/(15/2)=12/5 again.


So, we have two solutions of this problem.

Answer:

1: "\sqrt{\frac{10}{3}}," "\sqrt{\frac{15}{2}}," "\sqrt{\frac{6}{5}}," "\sqrt{\frac{24}{5}}" and the product is 12/5.

2: "\sqrt{\frac{5}{2}}," "\sqrt{10}," "\sqrt{\frac{8}{5}}," "\sqrt{\frac{18}{5}}" and the product is 12/5.